OpenStudy (anonymous):

simplify (2/9) two is to the 3rd power please help been stuck for 5 days :(

OpenStudy (anonymous):

so the term you want simplified is \[\frac{ 2^{3} }{ 9 }\] ok then, all you can do is calculate \[2^{3}\] which =2*2*2=8 so your answer is 8/9. you cannot make it and simpler than that

OpenStudy (anonymous):

thank you so much!

OpenStudy (isaiah.feynman):

You have just be saved!